Here is a problem that all my helpers say is not solvable, as do I.
Teacher says all problems are solvable
Three people own company shares. A and B together own 1/2, A and C own 1/3,
B and C own 1/5. What fraction do each own?
Can it be solved?
Thanks.

Answer by ankor@dixie-net.com(22740)

Three people own company shares.
A and B together own 1/2,
A and C own 1/3,
B and C own 1/5.
What fraction do each own?
:
I don't think this is possible, let say there are a total of 120 shares
then
A + B = 60
A + C = 40
B + C = 24
:
If A + B own 60, then C has to own the other 60
But it says that A + C = 40, we just decided that C has to own 60 ???